# Annealing Approximation in Master-Node Network Model

### v1

@@ Abstract
This paper investigates absorbing-state phase transitions in opinion dynamics through a master-node network model analyzed using annealing approximation. We develop a theoretical framework examining three fundamental regimes: systems converging to complete disagreement, complete consensus, or both states depending on initial conditions. The phase behavior is governed by two key chiral parameters - R measuring right-oriented influence and L measuring left-oriented influence in the network interactions. Our analysis reveals a rich phase diagram featuring both continuous and discontinuous transitions between disordered and ordered phases. The discontinuous transition emerges in systems with two absorbing states, where the final configuration depends critically on initial opinion distributions. The annealing approximation provides fundamental insights into how asymmetric social influences (chirality) shape collective opinion formation, acting as a symmetry-breaking element that drives the system toward polarization or consensus.
@@ 1 Introduction
These studies provide complementary perspectives on networked agent dynamics under master node influence. While [15] establishes the foundational framework and demonstrates phase transitions as key mechanisms in collective behavior, [16] offers detailed characterization of critical properties through exponent analysis. Together, they advance understanding of decentralized decision-making processes while contributing to non-equilibrium statistical mechanics. The current work builds upon these foundations by introducing a chiral master node formulation that incorporates geometric properties and Berry phase effects, enabling more comprehensive analysis of symmetry breaking patterns in opinion dynamics. The general formulation of the model, as shown in Section 2, contains so many parameters that it leads to a system that is not analytically accessible at first sight. Nevertheless, introducing an annealing approximation (on the variable associated with the influence of the master node) and examining the model at the mean-field level generates a system where several relevant cases can be characterized by just two quantities, as shown in Section 3. The analysis is then organized into three parts (Sections 4 to 6) and the role of the master node in this approximation is discussed in Section 7. Some final remarks can be found in Section 8.
@@ 2 General formulation (continuous approach)
The state of the above-defined model is completely described by the vector $s:=\begin{pmatrix}s_{1},&\cdots,&s_{N}\end{pmatrix}$, where $s_{i}\in\{0,1\}$ for each one of the $N$ vertices of the model. The dynamics is governed by the master equation
@@
which characterizes the time evolution of the probability $P$ of the system begin at state $s$ at time $t$. In (1), $s^{i}:=\begin{pmatrix}s_{1},&\cdots,&1-s_{i},&\cdots,&s_{N}\end{pmatrix}$ stands for the state where the $i$-th state is replaced by its complement and $w_{i}$ is the transition rate (there should be no confusion with the parameters $w_{1,2,3,4}$ introduced in the model) of the individual at site $i$.
@@
$s_{i-1}$ $s_{i+1}$ $0\rightarrow 0$ $0\rightarrow 1$ $1\rightarrow 0$ $1\rightarrow 1$
00 00 $p_{1}+\Gamma_{i}\left(q_{1}-p_{1}\right)$ $\overline{p_{1}}+\Gamma_{i}\left(\overline{q_{1}}-\overline{p_{1}}\right)$ $p_{2}+\Gamma_{i}\left(q_{2}-p_{2}\right)$ $\overline{p_{2}}+\Gamma_{i}\left(\overline{q_{2}}-\overline{p_{2}}\right)$
$1$ $1$ $p_{3}+\Gamma_{i}\left(q_{3}-p_{3}\right)$ $\overline{p_{3}}+\Gamma_{i}\left(\overline{q_{3}}-\overline{p_{3}}\right)$ $p_{4}+\Gamma_{i}\left(q_{4}-p_{4}\right)$ $\overline{p_{4}}+\Gamma_{i}\left(\overline{q_{4}}-\overline{p_{4}}\right)$
00 $1$ $w_{1}+\Gamma_{i}\left(v_{1}-w_{1}\right)$ $\overline{w_{1}}+\Gamma_{i}\left(\overline{v_{1}}-\overline{w_{1}}\right)$ $w_{2}+\Gamma_{i}\left(v_{2}-w_{2}\right)$ $\overline{w_{2}}+\Gamma_{i}\left(\overline{v_{2}}-\overline{w_{2}}\right)$
$1$ 00 $w_{3}+\Gamma_{i}\left(v_{3}-w_{3}\right)$ $\overline{w_{3}}+\Gamma_{i}\left(\overline{v_{3}}-\overline{w_{3}}\right)$ $w_{4}+\Gamma_{i}\left(v_{4}-w_{4}\right)$ $\overline{w_{4}}+\Gamma_{i}\left(\overline{v_{4}}-\overline{w_{4}}\right)$
Table 1: Transition rates for all possible neighbor configurations $(s_{i-1},s_{i+1})$ and state changes $(s_{i}\rightarrow s_{i}^{\prime})$. The rates depend on both the base parameters ($p_{j}$, $w_{j}$, $v_{j}$, $q_{j}$) and the master node influence $\Gamma_{i}$. The overline notation $\overline{x}$ denotes the complementary probability $1-x$.
@@
$w_{i}(s)$ $=s_{i}\big{\{}\left(1-s_{i-1}\right)\left(1-s_{i+1}\right)\left[p_{2}+\Gamma_{i}\left(q_{2}-p_{2}\right)\right]+s_{i-1}s_{i+1}\left[p_{4}+\Gamma_{i}\left(q_{4}-p_{4}\right)\right]+$
$+\left(1-s_{i-1}\right)s_{i+1}\left[w_{2}+\Gamma_{i}\left(v_{2}-w_{2}\right)\right]+s_{i-1}\left(1-s_{i+1}\right)\left[w_{4}+\Gamma_{i}\left(v_{4}-w_{4}\right)\right]\big{\}}+$
$+\left(1-s_{i}\right)\big{\{}\left(1-s_{i-1}\right)\left(1-s_{i+1}\right)\left[\overline{p_{1}}+\Gamma_{i}\left(\overline{q_{1}}-\overline{p_{1}}\right)\right]+s_{i-1}s_{i+1}\left[\overline{p_{3}}+\Gamma_{i}\left(\overline{q_{3}}-\overline{p_{3}}\right)\right]+$
@@
$=s_{i}\left[A_{i}^{(2)}+\left(B_{i}^{(4)}-A_{i}^{(2)}\right)s_{i-1}+\left(B_{i}^{(2)}-A_{i}^{(2)}\right)s_{i+1}+\left(A_{i}^{(2)}+A_{i}^{(4)}-B_{i}^{(2)}-B_{i}^{(4)}\right)s_{i-1}s_{i+1}\right]+$
@@
The magnetization of the $i$-th individual is defined as usual,
@@ 3 Annealed mean-field analysis
are introduced to investigate the dynamics (6) at the mean-field level. Furthermore, the annealed approach is assumed for the random variable $\Gamma_{i}$, which is then replaced by its expected value $\Gamma_{i}\rightarrow\langle\Gamma_{i}\rangle=\rho$, which is the fraction of connections from the master node. As a consequence of these hypothesis, all the subindices of $A_{i}^{(j)}\rightarrow A^{(j)}$, $\overline{A_{i}^{(j)}}\rightarrow\overline{A^{(j)}}$, $B_{i}^{(j)}\rightarrow B^{(j)}$ and $\overline{B_{i}^{(j)}}\rightarrow\overline{B^{(j)}}$ are now irrelevant and are going to be omitted. Furthermore, the dependence of $m(t)$ on time is also suppressed from now on in the notation.
@@
Here, the quantities $L$ and $R$ are defined as
@@
and play a major role in the analysis of the system.
At this point, it is interesting to restrict the analysis where some well-knowm absorbing states are present in the system, namely the stationary states $m^{\ast}=0$ and/or $m^{\ast}=1$ in the solution of (9). From (10), one can see that they correspond to the assumption $\overline{A^{(1)}}=0$ and $A^{(4)}=0$, respectively. In what follows, the analysis is divided into the following cases:
@@
4 System with absorbing state $m^{\ast}=0$ (without the absorbing state $m^{\ast}=1$) – Case I
In this scenario, one has $\overline{A^{(1)}}=0$, and the equation of motion is cast as $\dot{m}=F_{0}(m)$, where
@@ 4 System with absorbing state $m^{\ast}=0$ (without the absorbing state $m^{\ast}=1$) – Case I
If $A^{(4)}=0$, the system will also have $m^{\ast}=1$ as one of its stationary points. Since this case (00 and $1$ being both fixed points of the dynamical system) will be examined later, it will be assumed that $A^{(4)}\neq 0$ (or $A^{(4)}>0$) in this section. The stability analysis is going to be performed through the (one-dimensional) isocline method, which stems on the analysis of the sign of (12). Consequently, the relative positions of the points 00, $m_{+}^{(0)}$ and $m_{-}^{(0)}$ plays a major rule in this procedure and is detailed in the Appendix B, and the results are summarized in figure 1.
@@
with $\Omega_{0}^{(0)}\cup\Omega_{+}^{(0)}\cup\Omega_{0+}^{(0)}=\mathbb{R}^{2}$. As shown in figure 1, the regions $\Omega_{0}^{(0)}$ and $\Omega_{+}^{(0)}$ contain a single stationary point, which are $m^{\ast}=0$ and $m^{\ast}=m_{+}^{(0)}$, respectively. On the other hand, the region $\Omega_{0+}^{(0)}$ displays two stationary points, 00 and $m_{+}^{(0)}$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{0+}^{(0)}$. In Appendix B, it is shown that $10<m_{-}^{(0)}<m_{+}^{(0)}<$ 1 in this region, and
@@
In other words, the point $m=m_{-}^{(0)}$, which is an unstable fixed point, is a separatrix between these two stationary points, 00 and $m_{+}^{(0)}$.
@@
indicating a discontinuity of $m_{+}^{(0)}$ when crossing this boundary.
@@
5 System with absorbing state $m^{\ast}=1$ (without the absorbing state $m^{\ast}=0$) – Case II
In this scenario, one has $A^{(4)}=0$ and $\overline{A^{(1)}}=0$, and the equation of motion is cast as $\dot{m}=F_{1}(m)$, where
@@ 5 System with absorbing state $m^{\ast}=1$ (without the absorbing state $m^{\ast}=0$) – Case II
Analogously in the previous section, the case $\overline{A^{(1)}}=0$ is excluded in the analysis here, since it will be explored in the next section. Moreover, the technical details of the stationary analysis of 25 is shown in Appendix C, and the results can be summarized in the diagram of figure 2.
@@
with $\Omega_{1}^{(1)}\cup\Omega_{-}^{(1)}\cup\Omega_{1-}^{(1)}=\mathbb{R}^{2}$. As shown in figure 2, the regions $\Omega_{1}^{(1)}$ and $\Omega_{-}^{(1)}$ contain a single stationary point, which are $m^{\ast}=1$ and $m^{\ast}=m_{-}^{(1)}$, respectively. On the other hand, the region $\Omega_{1-}^{(1)}$ displays two stationary points, $1$ and $m_{-}^{(1)}$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{1-}^{(1)}$. In Appendix C, it is shown that $10<m_{-}^{(1)}<m_{+}^{(1)}<$ 1 in this region, and
@@
indicating a discontinuity of $m_{-}^{(1)}$ when crossing this boundary.
@@
6 System with absorbing state $m^{\ast}=0$ and $m^{\ast}=1$ – Case III
In this last scenario, one has $A^{(4)}=0$ and $\overline{A^{(1)}}=0$, and the equation of motion is cast as $\dot{m}=F_{01}(m)$, where

### v2

@@ Abstract
This paper investigates absorbing-state phase transitions in opinion dynamics through a master-node network model analyzed using annealing approximation. We develop a theoretical framework examining three fundamental regimes: systems converging to complete disagreement, complete consensus, or both states depending on initial conditions. The phase behavior is governed by two key chiral parameters: $R$ measuring right-oriented influence and $L$ measuring left-oriented influence in the network interactions. Our analysis reveals a rich phase diagram featuring both continuous and discontinuous transitions between disordered and ordered phases. The discontinuous transition emerges in systems with two absorbing states, where the final configuration depends critically on initial opinion distributions. The annealing approximation provides fundamental insights into how asymmetric social influences (chirality) shape collective opinion formation, acting as a symmetry-breaking element that drives the system toward polarization or consensus.
@@ 1 Introduction
These studies provide complementary perspectives on networked agent dynamics under master node influence. While [15] establishes the foundational framework and demonstrates phase transitions as key mechanisms in collective behavior, [16] offers detailed characterization of critical properties through exponent analysis. Together, they advance understanding of decentralized decision-making processes while contributing to nonequilibrium statistical mechanics. The current work builds upon these foundations by introducing a chiral master node formulation that incorporates geometric properties and Berry phase effects, enabling more comprehensive analysis of symmetry breaking patterns in opinion dynamics. The general formulation of the model, as shown in Section 2, contains so many parameters that it leads to a system that is not analytically accessible at first sight. Nevertheless, introducing an annealing approximation (on the variable associated with the influence of the master node) and examining the model at the mean-field level generates a system where several relevant cases can be characterized by just two quantities, as shown in Section 3. The analysis is then organized into three parts (Sections 4 to 6) and the role of the master node in this approximation is discussed in Section 7. Some final remarks can be found in Section 8.
@@ 2 General formulation (continuous approach)
The model consists of a one-dimensional chain where an individual (or spin) is assigned to each one of its $N$ sites. We denote by $s_{i}$ the state (or opinion) of individual $i$, who is located at position $i\in\{1,\ldots,N\}$ by construction. The periodic boundary condition is adopted ($s_{N+1}=s_{1}$) and the system can be seen as a “ring”, as shown in Figure 1. The spins have short-ranged interaction in the sense that an individual $i$ interacts with spins i1i-1 and $i+1$ only. A master node is then introduced to this setup and it can influence the way an individual interacts with its neighbors. However the master node is not affected by the state of the $N$ individuals of the ring.
Figure 1: Sketch of the system.
The state of the above-defined model is completely described by the vector $s:=(s_{1},\ldots,s_{N})$, where $s_{i}\in\{0,1\}$ for each one of the $N$ vertices of the model. The dynamics is governed by the master equation
@@
which characterizes the time evolution of the probability $P$ of the system begin at state $s$ at time $t$. In (1), $s^{i}:=(s_{1},\ldots,1-s_{i},\cdots,s_{N})$ denotes the state where the $i$th state is replaced by its complement and $w_{i}$ is the transition rate (there should be no confusion with the parameters $w_{1,2,3,4}$ introduced in the model) of the individual at site $i$.
@@
$s_{i-1}$ $s_{i+1}$ $0\rightarrow 0$ $0\rightarrow 1$ $\rightarrow 0 11$$\rightarrow 1$
0 0 $p_{1}+\Gamma_{i}\left(q_{1}-p_{1}\right)$ $\overline{p_{1}}+\Gamma_{i}\left(\overline{q_{1}}-\overline{p_{1}}\right)$ $p_{2}+\Gamma_{i}\left(q_{2}-p_{2}\right)$ $\overline{p_{2}}+\Gamma_{i}\left(\overline{q_{2}}-\overline{p_{2}}\right)$
11 11 $p_{3}+\Gamma_{i}\left(q_{3}-p_{3}\right)$ $\overline{p_{3}}+\Gamma_{i}\left(\overline{q_{3}}-\overline{p_{3}}\right)$ $p_{4}+\Gamma_{i}\left(q_{4}-p_{4}\right)$ $\overline{p_{4}}+\Gamma_{i}\left(\overline{q_{4}}-\overline{p_{4}}\right)$
0 11 $w_{1}+\Gamma_{i}\left(v_{1}-w_{1}\right)$ $\overline{w_{1}}+\Gamma_{i}\left(\overline{v_{1}}-\overline{w_{1}}\right)$ $w_{2}+\Gamma_{i}\left(v_{2}-w_{2}\right)$ $\overline{w_{2}}+\Gamma_{i}\left(\overline{v_{2}}-\overline{w_{2}}\right)$
11 0 $w_{3}+\Gamma_{i}\left(v_{3}-w_{3}\right)$ $\overline{w_{3}}+\Gamma_{i}\left(\overline{v_{3}}-\overline{w_{3}}\right)$ $w_{4}+\Gamma_{i}\left(v_{4}-w_{4}\right)$ $\overline{w_{4}}+\Gamma_{i}\left(\overline{v_{4}}-\overline{w_{4}}\right)$
Table 1: Transition rates for all possible neighbor configurations $(s_{i-1},s_{i+1})$ and state changes $(s_{i}\rightarrow s_{i}^{\prime})$. The rates depend on both the base parameters ($p_{j}$, $w_{j}$, $v_{j}$, $q_{j}$) and the master node influence $\Gamma_{i}$. The overline notation $\overline{x}$ denotes the complementary probability 1x1-x.
@@
$w_{i}(s)$ $=s_{i}\big\{\left(1-s_{i-1}\right)\left(1-s_{i+1}\right)\left[p_{2}+\Gamma_{i}\left(q_{2}-p_{2}\right)\right]+s_{i-1}s_{i+1}\left[p_{4}+\Gamma_{i}\left(q_{4}-p_{4}\right)\right]$
$+\left(1-s_{i-1}\right)s_{i+1}\left[w_{2}+\Gamma_{i}\left(v_{2}-w_{2}\right)\right]+s_{i-1}\left(1-s_{i+1}\right)\left[w_{4}+\Gamma_{i}\left(v_{4}-w_{4}\right)\right]\big\}$
$+\left(1-s_{i}\right)\big\{\left(1-s_{i-1}\right)\left(1-s_{i+1}\right)\left[\overline{p_{1}}+\Gamma_{i}\left(\overline{q_{1}}-\overline{p_{1}}\right)\right]+s_{i-1}s_{i+1}\left[\overline{p_{3}}+\Gamma_{i}\left(\overline{q_{3}}-\overline{p_{3}}\right)\right]$
@@
$=s_{i}\left[A_{i}^{(2)}+\left(B_{i}^{(4)}-A_{i}^{(2)}\right)s_{i-1}+\left(B_{i}^{(2)}-A_{i}^{(2)}\right)s_{i+1}+\left(A_{i}^{(2)}+A_{i}^{(4)}-B_{i}^{(2)}-B_{i}^{(4)}\right)s_{i-1}s_{i+1}\right]$
@@
The magnetization of the $i$th individual is defined as usual,
@@ 3 Annealed mean-field analysis
are introduced to investigate the dynamics (6) at the mean-field level. Furthermore, the annealed approach is assumed for the random variable $\Gamma_{i}$, which is then replaced by its expected value $\Gamma_{i}\rightarrow\langle\Gamma_{i}\rangle=\rho$, which is the fraction of connections from the master node. In this annealed approximation, the timescale of the fluctuation of the random variable $\Gamma_{i}$ is assumed to be much faster than the other variables of the system, which means that the influence of the master node over the individuals is felt by an effective one. From a social perspective, we can illustrate this scenario with the following example. Suppose that a central agent (such as mass media) controls the way the population interacts, and let the strength of this influence be represented by a real number in $0,1]\subset\mathbb{R}$. In the so-called quenched case, the central agent influences only a fraction $\rho$ of the population, but each affected individual experiences the full impact, with strength $1$. In the annealed scenario, the whole population is influenced, but the strength of the effect is just $\rho\in[0,1]$ for each individual.
As a consequence of these hypothesis (annealed and mean-field approach), all the subindices of $A_{i}^{(j)}\rightarrow A^{(j)}$, $\overline{A_{i}^{(j)}}\rightarrow\overline{A^{(j)}}$, $B_{i}^{(j)}\rightarrow B^{(j)}$ and $\overline{B_{i}^{(j)}}\rightarrow\overline{B^{(j)}}$ are now irrelevant and are going to be omitted. Furthermore, the dependence of $m(t)$ on time is also suppressed from now on in the notation.
@@
The term (10) can be regarded as the free-energy in dynamical Landau theory. Here, the quantities $L$ and $R$ are defined as
@@
and play a major role in the analysis of the system and can be interpreted in the following way. The quantities $R$ and $L$ emerge as fundamental parameters representing competing directional influences in the network. However, they do not correspond solely to movements towards one choice or another, as there are positive and negative terms in each of them. And yet, physically, a larger $R$ always corresponds to a right-oriented influence, combining the tendency to adopt right-neighbor opinions against resistance to such adoption. Conversely, $L$ quantifies left-oriented influence through similar mechanisms. Their ratio determines the system’s chiral character, with $R>L$ favoring right-biased consensus and vice versa. These parameters capture the inherent asymmetry in opinion propagation through the network. Interestingly, depending on those quantities, we observe a fixed point with coexistence of both opinions, like previous models with inflexibles [17, 18] or with asymmetric influences [19].
Note that, within this approach, the original model, which depends on more than 10 free parameters, can be cast as a function of only four ($R$, $L$, $\overline{A^{(1)}}$ and $A^{(4)}$), thereby rendering it amenable to analytical investigation. At this point, it is interesting to restrict the analysis where some well-known absorbing states are present in the system, namely the stationary states $m^{\ast}=0$ and/or $m^{\ast}=1$ in the solution of (9). From (10), one can see that they correspond to the assumption $\overline{A^{(1)}}=0$ and $A^{(4)}=0$, respectively. In what follows, the analysis is divided into the cases shown in Table 2.
@@
Table 2: Cases based on the absorbing states.
4 System with absorbing state $m^{\ast}=0$ (without the absorbing state $m^{\ast}=1$): Case I
In this scenario, one has $\overline{A^{(1)}}=0$, which means that the transition $000\rightarrow 010$ (see Table 1) is forbidden, and $A^{(4)}\neq 0$. The equation of motion is cast as $\dot{m}=F_{0}(m)$, where
@@ 4 System with absorbing state $m^{\ast}=0$ (without the absorbing state $m^{\ast}=1$): Case I
If $A^{(4)}=0$, the system will also have $m^{\ast}=1$ as one of its stationary points. Since this case (0 and $1$ being both fixed points of the dynamical system) will be examined later, it will be assumed that $A^{(4)}\neq 0$ (or $A^{(4)}>0$) in this section. The stability analysis is going to be performed through the (one-dimensional) isocline method, which stems on the analysis of the sign of (12). Consequently, the relative positions of the points 0, $m_{+}^{(0)}$ and $m_{-}^{(0)}$ plays a major rule in this procedure and is detailed in the Appendix A, and the results are summarized in figure 2.
@@
with $\Omega_{0}^{(0)}\cup\Omega_{+}^{(0)}\cup\Omega_{0+}^{(0)}=\mathbb{R}^{2}$. As shown in figure 2, the regions $\Omega_{0}^{(0)}$ and $\Omega_{+}^{(0)}$ contain a single stationary point, which are $m^{\ast}=0$ and $m^{\ast}=m_{+}^{(0)}$, respectively. On the other hand, the region $\Omega_{0+}^{(0)}$ displays two stationary points, 0 and $m_{+}^{(0)}$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{0+}^{(0)}$. In Appendix A, it is shown that $0<m_{-}^{(0)}<m_{+}^{(0)}<1$ in this region, and
@@
In other words, the point $m=m_{-}^{(0)}$, which is an unstable fixed point, is a separatrix between these two stationary points, 0 and $m_{+}^{(0)}$.
@@
indicating a discontinuity of $m_{+}^{(0)}$ when crossing this boundary. More precisely, the underlying mechanism of this discontinuity is a saddle-node bifurcation [20]: crossing from the interior of $\Omega_{0}^{(0)}$ to the interior of $\Omega_{0+}^{(0)}$, a stationary point different from $m^{\ast}=0$, given by (24), emerges in the boundary $R=\sqrt{4LA^{(4)}}$; this point bifurcates into $m_{+}^{(0)}$ (stable) and $m_{-}^{(0)}$ (unstable) in $\Omega_{0+}^{(0)}$.
@@
5 System with absorbing state $m^{\ast}=1$ (without the absorbing state $m^{\ast}=0$): Case II
In this scenario, one has $A^{(4)}=0$, which forbids the transition $111\rightarrow 101$ (see Table 1), and $\overline{A^{(1)}}\neq 0$. The equation of motion is cast as $\dot{m}=F_{1}(m)$, where
@@ 5 System with absorbing state $m^{\ast}=1$ (without the absorbing state $m^{\ast}=0$): Case II
Analogously in the previous section, the case $\overline{A^{(1)}}=0$ is excluded in the analysis here, since it will be explored in the next section. Moreover, the technical details of the stationary analysis of 25 is shown in Appendix B, and the results can be summarized in the diagram of figure 3.
@@
with $\Omega_{1}^{(1)}\cup\Omega_{-}^{(1)}\cup\Omega_{1-}^{(1)}=\mathbb{R}^{2}$. As shown in figure 3, the regions $\Omega_{1}^{(1)}$ and $\Omega_{-}^{(1)}$ contain a single stationary point, which are $m^{\ast}=1$ and $m^{\ast}=m_{-}^{(1)}$, respectively. On the other hand, the region $\Omega_{1-}^{(1)}$ displays two stationary points, $1$ and $m_{-}^{(1)}$; the initial condition $m(t=0)=m_{0}$ determines the trajectory toward one of these two fixed points in $\Omega_{1-}^{(1)}$. In Appendix B, it is shown that $0<m_{-}^{(1)}<m_{+}^{(1)}<1$ in this region, and
@@
indicating a discontinuity of $m_{-}^{(1)}$ when crossing this boundary. More precisely, the underlying mechanism of this discontinuity is a saddle-node bifurcation [20]: Crossing from the interior of $\Omega_{1}^{(1)}$ to the interior of $\Omega_{1-}^{(1)}$, a stationary point different from $m^{\ast}=1$, given by (37), emerges in the boundary $L=\sqrt{4R\overline{A^{(1)}}}$; this point bifurcates into $m_{-}^{(1)}$ (stable) and $m_{+}^{(1)}$ (unstable) in $\Omega_{1-}^{(1)}$.
@@
6 System with absorbing state $m^{\ast}=0$ and $m^{\ast}=1$: Case III
In this last scenario, one has $A^{(4)}=0$ and $\overline{A^{(1)}}=0$, which forbid the transitions $111\rightarrow 101$ and $000\rightarrow 010$ (see Table 1). The equation of motion is cast as $\dot{m}=F_{01}(m)$, where